# RocketKV: Accelerating Long-Context LLM Inference via Two-Stage KV Cache Compression

### v1

@@ 5 Conclusion
RocketKV presents a novel, training-free approach to KV cache compression, addressing the challenges of memory bandwidth and capacity demands during the decode phase of LLM inference. RocketKV contains two consecutive stages: SnapKV++ for coarse-grain KV cache eviction and hybrid attention for fine-grain dynamic KV token selection. Our evaluations on models and long-context benchmarks demonstrate that RocketKV maintains comparable accuracy to full KV cache attention while significantly lowering memory bandwidth and capacity usage. At low token budgets of 256 or 512, RocketKV outperforms all other techniques by a considerable margin. It achieves up to 3$\times$ speedup and 31% peak memory reduction at the decode phase on an NVIDIA H100 GPU compared to the full KV cache baseline, highlighting its efficiency and potential for widespread application in optimizing LLM performance.
@@ A.1 Models
We evaluate our methods on the three widely used long-context models: Llama3.1-8B-Ins [1], Mistral-7B-Ins-v0.2 [13], and LongChat-7B-v1.5 [4]. They are all decoder-only transformer models [33]. Llama3.1-8B-Ins and Mistral-7B-Ins-v0.2 use GQA, while LongChat-7B-v1.5 uses MHA in the attention module. In terms of sequence length, Llama3.1-8B-Ins supports a maximum sequence length of 128K while the other two support up to 32K sequence length. Prior work [31, 6] usually skips the first two layers during KV cache compression to maintain accuracy. We found this is only necessary for LongChat-7B-v1.5; so, we conduct KV cache compression on all attention layers for Llama3.1-8B-Ins and Mistral-7B-Ins-v0.2 but skip the first two layers for LongChat-7B-v1.5.
@@ A.2 Benchmarks
Our LongBench and Needle-in-a-Haystack settings are mostly following the prior work [34]. For LongBench, we set the maximum prompt length to 127,500 for Llama3.1-8B-Ins and 31,500 for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5. Prompts beyond the maximum prompt length will be middle-truncated by keeping the first and last half of the input tokens. For Needle-in-a-Haystack, we evaluate with 10 different input sequence lengths uniformly spanning from 2048 to 81,920 words and 10 different depths for each sequence length in Llama3.1-8B-Ins. The maximum word count of 81,920 roughly converts to 109K tokens by the tokenizer in Llama3.1-8B-Ins. For the other two models, the input sequence lengths range from 512 to 20480 words since they support up to 32K sequence length, which convert up to  30K and 31K tokens for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5, respectively. For RULER, we mostly follow the configurations in the original benchmark except for reducing the number of examples per task from 500 to 50 to speed up the evaluation.
@@ A.3 Baselines
Since RocketKV primarily focuses on accelerating the decode phase of LLM inference, we do not perform any KV cache compression or sparse attention mechanism in the prefill phase across all comparing baselines including DuoAttention for fair comparison.
@@
DuoAttention: We set the initial and recent tokens to 128 and 512 respectively for streaming heads if the token budget is larger than 640, and vary the ratio of retrieval heads to match with the given token budget on average. If the token budget is lower than 640, all attention heads become streaming heads. In this case, we set the number of initial tokens for attention sink to 128 if the token budget is larger than 256, otherwise it is set to 20% of the token budget. The rest of the token budget is assigned to recent tokens within the sliding window.
SnapKV: We set the observation window size to 32 and kernel size for pooling to 7 following the SnapKV paper. Since SnapKV only compresses the KV cache of the input prompt, the token budget for the input prompt is calculated by deducting the number of maximum generated tokens from the total token budget. This budget is further divided by the number of heads per attention group in case of GQA as SnapKV performs on each attention head separately and the pruned KV cache is not shared within the attention group.
Quest: We evenly split the token budget between approximate attention for identifying top-k KV token indices and top-k sparse attention so that the token budget can accurately reflect total memory fetching bandwidth required in the attention module. The original Quest method is not compatible with GQA. We modify it to select indices of top-k accumulated attention scores per attention group rather than per attention head similar to our enhancement in SnapKV++.
@@
RocketKV: We evenly split the compression ratio between SnapKV++ in the first stage and hybrid attention in the second stage. For SnapKV++, we use an observation window size of 32 and adaptively select the kernel size for pooling between 63 and 511, depending on the input sequence length. For Llama3.1-8B-Ins, we use the kernel size of 63 if the input sequence length is shorter than 48K, otherwise 511 is used. For LongChat-7B-v1.5, the input sequence length threshold is 12K, and Mistral-7B-Ins-v0.2 always selects a kernel size of 63. The adaptive selection policy is empirically determined based on our ablation study in Section 4.4. For hybrid attention, we evenly split the token budget between approximate attention for identifying top-k indices and top-k sparse attention.
Appendix B Detailed Results
B.1 Further Ablation Study
This subsection presents additional results from the ablation studies discussed in the main paper. These results further enhance our observations in Section 4.4.
(a) Llama3.1-8B-Ins, SeqLen= 16K
(b) Llama3.1-8B-Ins, SeqLen=32K
(c) Llama3.1-8B-Ins, SeqLen=64K
(d) Llama3.1-8B-Ins, SeqLen=96K
(e) Mistral-7B-Ins-v0.2, SeqLen=8K
(f) Mistral-7B-Ins-v0.2, SeqLen=16K
(g) Mistral-7B-Ins-v0.2, SeqLen=24K
(h) Mistral-7B-Ins-v0.2, SeqLen=32K
Figure 9: RULER accuracy comparison between SnapKV and SnapKV with GQA enhancement.
(a) LongBench, Llama3.1-8B-Ins
(b) LongBench, Mistral-7B-Ins-v0.2
(c) LongBench, LongChat-7B-v1.5
(d) NIAH, Llama3.1-8B-Ins
(e) NIAH, Mistral-7B-Ins-v0.2
(f) NIAH, LongChat-7B-v1.5
Figure 10: Comparing RocketKV with different kernel sizes (7, 63 and 511) under LongBench and Needle-in-a-Haystack
(a) Llama3.1-8B-Ins, SeqLen = 16K
(b) Llama3.1-8B-Ins, SeqLen= 32K
(c) Llama3.1-8B-Ins, SeqLen=64K
(d) Llama3.1-8B-Ins, SeqLen=96K
(e) Mistral-7B-Ins-v0.2, SeqLen=8K
(f) Mistral-7B-Ins-v0.2, SeqLen=16K
(g) Mistral-7B-Ins-v0.2, SeqLen=24K
(h) Mistral-7B-Ins-v0.2, SeqLen=32K
(i) LongChat-7B-v1.5, SeqLen=8K
(j) LongChat-7B-v1.5, SeqLen=16K
(k) LongChat-7B-v1.5, SeqLen=24K
(l) LongChat-7B-v1.5, SeqLen=32K
Figure 11: accuracy comparison among Hybrid, Quest, and SparQ under RULER with various sequence lengths
@@ B.2 Needle-in-a-Haystack Visualization
Needle-in-a-haystack (NIAH) is a type of synthetic challenge designed to test how effectively an LLM can retrieve specific information in a large volume of text [16]. In Figures 1213,  14, the x-axis shows the document length (i.e., ”haystack”), while the y-axis marks the relative position of the ”needle” (i.e., a short sentence) within the text. As the results show, RocketKV can retrieve the needle with almost the same accuracy as the Full-KV throughout the whole text across all three models with token budgets as low as 512, and only incurs small accuracy losses with a token budget of 256.
@@ B.3 Detailed Accuracy Results
In this section, we present 18 detailed tables that display the accuracy results discussed in previous sections. These tables offer additional information and insights into the results we captured.
@@
Hybrid 31.2 45.7 54.7 54.3 46.4 31.6 34.9 25.3 27.2 72.5 91.5 44.2 99.5 63.7 58.6 7.5
RocketKV 30.4 45.3 54.9 54.2 46.4 30.5 34.2 25.3 27.3 74.0 91.5 43.7 99.5 65.7 60.2 7.5
Table 4: Results of RULER, SeqLen=16000 for Llama3.1-8B-Ins
@@
Hybrid 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 77.0 91.3 82.0 50.0 92.3
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 59.6 94.7 82.0 50.0 91.2
Table 5: Results of RULER, SeqLen=32000 for Llama3.1-8B-Ins
@@
Hybrid 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.0 93.2 19.8 79.3 74.0 52.0 85.9
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.0 93.6 15.8 90.0 76.0 52.0 86.6
Table 6: Results of RULER, SeqLen=64000 for Llama3.1-8B-Ins
@@
Hybrid 100.0 100.0 100.0 100.0 98.0 96.0 96.5 98.0 90.8 0.2 70.0 76.0 56.0 83.2
RocketKV 100.0 100.0 100.0 100.0 98.0 100.0 98.0 99.5 92.4 0.0 84.0 74.0 56.0 84.8
Table 7: Results of RULER, SeqLen=96000 for Llama3.1-8B-Ins
@@
Hybrid 27.8 32.2 49.0 42.0 27.4 19.2 33.0 23.9 27.2 71.5 86.2 42.7 83.3 57.9 56.2 2.3
RocketKV 27.7 32.1 48.7 41.2 27.8 19.6 32.2 24.6 27.0 71.5 86.2 42.0 82.2 57.8 56.2 3.1
Table 10: Results of RULER, SeqLen=8000 for Mistral-7B-Ins-v0.2
@@
Hybrid 100.0 100.0 100.0 98.0 100.0 82.0 79.5 96.0 97.6 90.6 74.0 80.0 64.0 89.4
RocketKV 100.0 100.0 94.0 98.0 100.0 82.0 91.0 98.0 98.8 87.2 74.0 84.0 66.0 90.2
Table 11: Results of RULER, SeqLen=16000 for Mistral-7B-Ins-v0.2
@@
Hybrid 100.0 100.0 100.0 84.0 92.0 78.0 84.0 86.5 84.4 74.6 81.3 78.0 58.0 84.7
RocketKV 100.0 100.0 100.0 82.0 88.0 70.0 95.5 98.0 88.8 70.0 86.0 74.0 56.0 85.3
Table 12: Results of RULER, SeqLen=24000 for Mistral-7B-Ins-v0.2
@@
Hybrid 100.0 98.0 90.0 88.0 92.0 78.0 73.5 83.0 94.8 62.0 82.7 70.0 58.0 82.3
RocketKV 100.0 100.0 98.0 82.0 92.0 70.0 87.5 94.0 97.2 47.8 85.3 60.0 54.0 82.1
Table 13: Results of RULER, SeqLen=32000 for Mistral-7B-Ins-v0.2
@@
Hybrid 20.8 29.2 43.2 33.5 24.2 14.1 31.0 22.9 26.5 66.0 84.2 12.9 28.5 55.3 59.3 0.0
RocketKV 20.1 29.8 43.2 33.2 25.1 14.4 30.2 22.7 26.6 67.0 84.3 12.8 28.5 54.9 59.4 0.0
Table 16: Results of RULER, SeqLen=8000 for LongChat-7B-v1.5
@@
Hybrid 100.0 100.0 96.0 98.0 100.0 80.0 96.5 69.5 92.8 34.4 82.7 66.0 58.0 82.6
RocketKV 100.0 100.0 96.0 98.0 100.0 74.0 96.5 71.5 92.8 33.6 82.0 66.0 58.0 82.2
Table 17: Results of RULER, SeqLen=16000 for LongChat-7B-v1.5
@@
Hybrid 100.0 100.0 96.0 92.0 98.0 52.0 83.0 53.0 79.2 26.4 70.7 62.0 44.0 73.6
RocketKV 100.0 100.0 96.0 92.0 96.0 40.0 82.0 51.5 79.2 18.8 73.3 62.0 42.0 71.8
Table 18: Results of RULER, SeqLen=24000 for LongChat-7B-v1.5
@@
Hybrid 100.0 100.0 98.0 88.0 82.0 22.0 58.5 38.5 67.2 12.2 66.0 46.0 34.0 62.5
RocketKV 100.0 100.0 98.0 92.0 88.0 16.0 65.0 39.0 72.8 3.6 68.7 50.0 32.0 63.5
Table 19: Results of RULER, SeqLen=32000 for LongChat-7B-v1.5

### v2

@@ 5 Conclusion
RocketKV presents a novel, training-free approach to KV cache compression, addressing the challenges of memory bandwidth and capacity demands during the decode phase of LLM inference. RocketKV contains two consecutive stages: SnapKV for coarse-grain permanent KV cache eviction and hybrid sparse attention (HSA) for fine-grain dynamic KV token selection. Our evaluations on various models and long-context benchmarks demonstrate that RocketKV maintains comparable accuracy to full KV cache attention while significantly lowering memory bandwidth and capacity usage with a compression ratio of up to 400$\times$, as well as up to 3.7$\times$ end to end speedup and 32.6% peak memory reduction at the decode phase, highlighting its efficiency and potential for widespread application in optimizing LLM performance. We also propose a variant of RocketKV called RocketKV-MT that is optimized for multi-turn scenarios.
Acknowledgment
We thank the anonymous ICML reviewers for their valuable and constructive feedback. We are grateful to David Nellans for overseeing the internship process of Payman Behnam, during which this work was supported by the NVIDIA Graduate Fellowship. We also thank Song Han and Christos Kozyrakis for their insightful discussions and helpful feedback.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Models
We evaluate our methods on the three widely used long-context models: Llama3.1-8B-Ins (AI@Meta, 2024), Mistral-7B-Ins-v0.2 (Jiang et al., 2023), and LongChat-7B-v1.5 (Bai et al., 2024). They are all decoder-only Transformer models (Vaswani et al., 2017). Llama3.1-8B-Ins and Mistral-7B-Ins-v0.2 use GQA, while LongChat-7B-v1.5 uses MHA in the attention module. In terms of sequence length, Llama3.1-8B-Ins supports a maximum sequence length of 128K while the other two support up to 32K sequence length. Prior work (Tang et al., 2024; Chen et al., 2025) usually skips the first two layers during KV cache compression to maintain accuracy. We found that this is only necessary for LongChat-7B-v1.5; so, we conduct KV cache compression on all attention layers for Llama3.1-8B-Ins and Mistral-7B-Ins-v0.2 but skip the first two layers for LongChat-7B-v1.5.
@@ A.2 Benchmarks
Our LongBench and Needle-in-a-Haystack settings mostly follow the prior work (Yuan et al., 2024). For LongBench, we set the maximum prompt length to 127,500 for Llama3.1-8B-Ins and 31,500 for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5. Prompts beyond the maximum prompt length will be middle-truncated by keeping the first and last half of the input tokens. For Needle-in-a-Haystack, we evaluate with 10 different input sequence lengths uniformly spanning from 2048 to 81,920 words and 10 different depths for each sequence length in Llama3.1-8B-Ins. The maximum word count of 81,920 roughly converts to 109K tokens by the tokenizer in Llama3.1-8B-Ins. For the other two models, the input sequence lengths range from 512 to 20480 words since they support up to 32K sequence length, which convert up to about 30K and 31K tokens for Mistral-7B-Ins-v0.2 and LongChat-7B-v1.5, respectively. For RULER, we mostly follow the configurations in the original benchmark, except for reducing the number of examples per task from 500 to 50 to speed up the evaluation. For SCBench, we follow the multi-turn setting and set the maximum prompt length to 127,500 for Llama3.1-8B-Ins.
@@ A.3 Baselines
Since RocketKV primarily focuses on accelerating the decode phase of LLM inference, we do not perform any KV cache compression or sparse attention mechanism in the prefill phase across all comparing baselines, including DuoAttention, for a fair comparison.
@@
DuoAttention: We set the initial and recent tokens to 128 and 512, respectively, for streaming heads if the token budget is larger than 640, and vary the ratio of retrieval heads to match the given token budget on average. If the token budget is lower than 640, all attention heads become streaming heads. In this case, we set the number of initial tokens for the attention sink to 128 if the token budget is larger than 256, otherwise, it is set to 20% of the token budget. The rest of the token budget is assigned to recent tokens within the sliding window.
SnapKV: We set the observation window size to 32 for all single-turn benchmarks and 128 for multi-turn SCBench. The kernel size for pooling is set to 7, following the SnapKV paper. The token budget is divided by the number of heads per attention group in case of GQA as the SnapKV performs on each attention head separately and the pruned KV cache is not shared within the attention group.
Quest: We evenly split the token budget between approximate attention for identifying top-k KV token indices and top-k sparse attention so that the token budget can accurately reflect the total memory fetching bandwidth required in the attention module. The original Quest method is not compatible with GQA. We modify it to select indices of top-k accumulated attention scores per attention group rather than per attention head.
@@
RocketKV/RocketKV-MT: For SnapKV in the first stage, We set the observation window size to 32 for all single-turn benchmarks and 128 for multi-turn SCBench. For HSA in the second stage, we evenly split the token budget between approximate attention for identifying top-k indices and top-k sparse attention.
Appendix B Additional Results
B.1 Ablation Studies
In this subsection, we present a series of ablation studies to further demonstrate the effectiveness of RocketKV.
(a) LongBench
(b) NIAH
(c) RULER, SeqLen = 16K
(d) RULER, SeqLen= 32K
(e) RULER, SeqLen=64K
(f) RULER, SeqLen=96K
Figure 8: Accuracy comparison among HSA, Quest, and SparQ on Llama3.1-8B-Ins.
(a) LongBench
(b) NIAH
(c) RULER, SeqLen=16K
(d) RULER, SeqLen=32K
(e) RULER, SeqLen=64K
(f) RULER, SeqLen=96K
Figure 9: Comparing adaptive against static split factors on Llama3.1-8B-Ins.
B.1.1 Comparing HSA, Quest, and SparQ
To illustrate the effectiveness of hybrid sparse attention (HSA), we compare the accuracy of the standalone HSA mechanism against Quest and SparQ. Figure 8 demonstrates the results on Llama3.1-8B-Ins across multiple different benchmarks. In all cases, HSA consistently outperforms Quest and SparQ, especially at low token budgets. This clearly demonstrates the advantage of HSA, which intelligently leverages approximations in both sequence and head dimensions compared to single dimension approximation methods such as Quest and SparQ.
B.1.2 Split Factor
In this study, we compare the adaptive compression decomposition method against statically determined split factors $r$ ranging from 0.3 to 0.7. A split factor of 0.5 indicates an even split between the first and second stage of RocketKV. As shown in Figure 9, the best static split factor varies with different sequence lengths and token budgets, while adaptive compression decomposition provides comparable accuracy to the best static split factor in most cases.
@@ B.2 Needle-in-a-Haystack Visualization
Needle-in-a-haystack (NIAH) is a type of synthetic challenge designed to test how effectively an LLM can retrieve specific information in a large volume of text (Kamradt, 2023). In Figures 1011,  12, the x-axis shows the document length (i.e., “haystack”), while the y-axis marks the relative position of the “needle” (i.e., a short sentence) within the text. As shown in the results, RocketKV can retrieve the needle with almost the same accuracy as Full-KV throughout the whole text across all three models with token budgets as low as 256.
@@ B.3 Detailed Accuracy Results
In this section, we present 20 detailed tables that display the accuracy results discussed in previous sections. These tables offer additional information and insights into the results we captured.
@@
HSA 31.2 45.7 54.7 54.3 46.4 31.6 34.9 25.3 27.2 72.5 91.5 44.2 99.5 63.7 58.6 7.5
RocketKV 30.3 45.3 54.3 54.8 46.3 30.4 34.7 25.0 27.1 73.0 91.5 44.3 99.5 63.7 58.5 7.6
Table 4: Average SCBench results for Llama3.1-8B-Ins
Token Budget
Method
Retr.String
Retr.Semantic
Global
Multi-task
Avg.
N/A Full-KV 49.5 40.7 36.2 64.6 49.9
1024 Exact-TopK 32.3 41.0 36.8 61.8 45.0
DuoAttention 0.1 26.6 33.5 11.6 20.9
SnapKV 0.2 17.0 34.7 10.7 17.6
Quest 0.0 22.4 25.2 4.0 15.7
SparQ 0.0 19.2 23.6 0.9 13.6
HSA 4.1 27.2 27.2 20.0 22.1
RocketKV 14.3 22.7 35.3 26.2 26.6
RocketKV-MT 40.5 35.7 36.8 56.2 44.3
2048 Exact-TopK 38.0 41.2 36.6 62.9 46.7
DuoAttention 0.1 26.6 33.9 12.2 21.2
SnapKV 1.1 17.1 35.9 12.1 18.5
Quest 2.0 24.4 26.6 11.8 18.8
SparQ 0.1 21.1 25.8 4.1 15.5
HSA 8.4 35.3 35.7 53.4 34.6
RocketKV 16.5 25.0 35.4 30.5 28.8
RocketKV-MT 42.2 38.4 37.0 59.5 46.4
4096 Exact-TopK 43.4 42.1 36.4 63.8 48.7
DuoAttention 0.1 26.7 33.9 12.7 21.3
SnapKV 3.5 19.1 35.1 13.2 19.9
Quest 8.6 26.8 28.5 23.0 24.0
SparQ 3.0 29.9 29.2 30.8 25.2
HSA 13.1 38.7 36.3 60.9 38.6
RocketKV 18.9 28.5 35.9 33.9 31.5
RocketKV-MT 44.8 39.9 36.8 60.3 47.8
8192 Exact-TopK 48.0 42.4 36.4 64.7 50.2
DuoAttention 0.1 27.3 34.0 16.5 22.2
SnapKV 5.0 20.3 35.7 19.0 21.9
Quest 10.1 31.6 32.0 31.9 28.8
SparQ 6.8 41.5 35.6 64.1 38.3
HSA 34.1 41.8 36.5 63.8 46.0
RocketKV 22.9 33.2 35.8 39.1 35.2
RocketKV-MT 47.7 42.0 36.2 62.9 49.6
16384 Exact-TopK 49.5 42.8 36.4 65.0 50.8
DuoAttention 0.4 28.9 34.0 30.5 25.4
SnapKV 6.9 22.0 34.5 24.7 23.8
Quest 12.5 40.6 36.2 61.8 39.3
SparQ 40.3 42.4 36.1 64.5 48.0
HSA 46.5 43.1 36.6 64.4 50.0
RocketKV 29.4 37.2 36.1 44.6 39.5
RocketKV-MT 49.9 42.1 36.3 64.3 50.5
Table 5: Individual SCBench results for Llama3.1-8B-Ins
Token Budget
Method
kv
prefix.suffix
vt
repoqa
qa.eng
qa.chn
choice.eng
many.shot
summary
mf
summary.needles
repoqa.and.kv
N/A Full-KV 76.2 34.6 37.6 48.0 21.5 20.9 72.6 59.3 37.5 12.0 60.4 68.8
1024 Exact-TopK 43.0 18.6 35.4 42.0 30.1 19.3 72.5 60.4 36.5 13.7 59.3 64.2
DuoAttention 0.2 0.0 0.1 0.5 18.3 15.8 71.7 60.0 28.3 12.2 22.9 0.3
SnapKV 0.0 0.0 0.6 0.9 8.7 12.2 46.2 69.6 27.2 7.2 21.2 0.1
Quest 0.0 0.0 0.1 0.0 7.2 10.1 72.1 50.7 11.7 13.2 8.0 0.0
SparQ 0.0 0.0 0.0 0.0 4.6 3.7 68.6 47.8 6.5 16.5 1.9 0.0
HSA 0.0 0.0 12.4 5.7 22.9 9.0 71.2 59.6 9.7 12.2 35.3 4.7
RocketKV 10.2 2.4 30.4 5.0 16.5 15.2 54.1 65.6 30.0 10.3 49.1 3.3
RocketKV-MT 64.4 21.8 35.2 20.5 31.3 18.5 72.5 61.5 36.2 12.7 61.1 51.4
2048 Exact-TopK 53.4 23.8 36.7 43.9 29.6 18.7 72.5 60.0 37.2 12.5 59.7 66.1
DuoAttention 0.2 0.0 0.1 0.5 18.8 16.0 71.2 61.1 28.9 11.8 24.0 0.3
SnapKV 0.0 0.2 3.2 1.4 9.9 12.0 44.9 70.0 27.9 9.7 23.6 0.6
Quest 0.0 0.0 5.9 0.0 12.8 12.8 72.1 50.0 14.1 15.5 23.2 0.3
SparQ 0.0 0.0 0.4 0.0 9.3 6.9 68.1 61.1 6.0 10.2 8.2 0.0
HSA 0.0 0.0 25.3 23.6 29.6 16.4 71.7 61.1 31.9 14.0 57.2 49.6
RocketKV 11.6 4.2 33.7 8.9 19.7 14.8 56.8 64.1 32.3 9.8 55.1 6.0
RocketKV-MT 61.2 30.6 34.7 31.8 30.4 19.5 71.7 61.5 36.9 12.5 61.0 58.1
4096 Exact-TopK 66.0 26.8 37.4 46.6 29.6 18.3 73.9 59.6 37.5 12.2 59.7 67.9
DuoAttention 0.2 0.0 0.0 0.5 19.7 15.4 71.2 60.7 28.6 12.5 25.1 0.3
SnapKV 0.0 0.2 10.3 2.7 12.2 12.7 48.9 67.8 28.6 8.8 25.6 0.9
Quest 0.0 0.0 25.7 0.5 21.4 13.5 71.7 55.9 16.5 13.0 44.8 1.3
SparQ 0.0 0.0 9.0 17.0 22.3 8.8 71.2 60.0 16.2 11.3 42.6 18.9
HSA 4.0 0.8 34.4 34.3 29.3 18.6 72.5 59.3 36.4 13.3 60.3 61.5
RocketKV 17.6 4.8 34.2 13.2 20.3 17.8 62.9 62.2 34.6 10.8 58.2 9.7
RocketKV-MT 65.8 32.6 36.1 37.5 30.7 18.5 73.0 60.7 37.4 12.3 60.4 60.2
8192 Exact-TopK 72.6 33.0 38.3 45.9 30.7 18.9 73.9 59.6 37.4 12.2 60.4 68.9
DuoAttention 0.2 0.0 0.0 0.5 21.2 16.0 71.7 60.7 29.3 11.8 32.4 0.6
SnapKV 0.0 0.2 14.9 4.1 13.0 14.4 49.7 69.6 28.5 9.0 36.5 1.4
Quest 0.2 0.0 30.0 10.7 27.5 17.1 71.2 59.6 22.3 14.0 44.3 19.6
SparQ 0.0 0.4 20.0 50.7 27.8 15.6 71.7 59.3 35.5 12.2 60.5 67.6
HSA 47.4 20.0 34.8 44.8 29.9 19.7 73.0 59.6 36.8 13.0 60.1 67.5
RocketKV 25.0 7.4 36.3 23.2 23.5 17.8 68.2 60.4 36.2 10.7 60.6 17.6
RocketKV-MT 70.4 35.0 37.7 44.3 31.0 18.7 73.9 60.0 37.2 11.5 60.5 65.2
16384 Exact-TopK 74.8 35.2 38.5 48.4 30.7 19.1 73.0 59.3 37.4 12.3 60.4 69.5
DuoAttention 0.2 0.0 1.1 3.6 24.1 16.7 71.2 59.6 29.9 12.5 46.4 14.5
SnapKV 0.0 0.4 20.4 5.9 14.2 14.5 53.2 64.1 29.9 9.5 47.8 1.7
Quest 2.8 2.2 32.6 43.0 28.7 18.2 72.6 59.6 36.1 13.0 59.8 63.9
SparQ 60.8 26.0 34.1 48.4 29.9 18.0 73.4 59.3 36.6 12.3 60.5 68.5
HSA 70.8 30.8 38.0 50.0 29.9 19.0 73.4 59.6 37.6 12.7 60.7 68.2
RocketKV 39.6 10.8 37.6 33.0 25.5 18.7 71.7 58.9 37.3 12.2 60.9 28.3
RocketKV-MT 74.6 37.2 38.0 46.1 30.1 18.9 73.4 59.6 37.6 11.7 60.8 67.8
Table 6: Results of RULER-16K for Llama3.1-8B-Ins
@@
HSA 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 77.0 91.3 82.0 50.0 92.3
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.2 72.8 93.3 80.0 50.0 91.9
Table 7: Results of RULER-32K for Llama3.1-8B-Ins
@@
HSA 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.0 93.2 19.8 79.3 74.0 52.0 85.9
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.0 93.6 17.2 87.3 76.0 52.0 86.5
Table 8: Results of RULER-64K for Llama3.1-8B-Ins
@@
HSA 100.0 100.0 100.0 100.0 98.0 96.0 96.5 98.0 90.8 0.2 70.0 76.0 56.0 83.2
RocketKV 100.0 100.0 100.0 100.0 100.0 100.0 96.5 100.0 92.8 1.2 76.0 76.0 56.0 84.5
Table 9: Results of RULER-96K for Llama3.1-8B-Ins
@@
HSA 27.8 32.2 49.0 42.0 27.4 19.2 33.0 23.9 27.2 71.5 86.2 42.7 83.3 57.9 56.2 2.3
RocketKV 27.8 32.1 48.7 41.8 27.5 19.2 33.4 24.5 27.1 71.5 86.2 42.5 82.8 56.0 54.0 2.9
Table 12: Results of RULER-8K for Mistral-7B-Ins-v0.2
@@
HSA 100.0 100.0 100.0 98.0 100.0 82.0 79.5 96.0 97.6 90.6 74.0 80.0 64.0 89.4
RocketKV 100.0 100.0 94.0 98.0 100.0 82.0 91.0 98.0 99.2 90.4 73.3 84.0 66.0 90.5
Table 13: Results of RULER-16K for Mistral-7B-Ins-v0.2
@@
HSA 100.0 100.0 100.0 84.0 92.0 78.0 84.0 86.5 84.4 74.6 81.3 78.0 58.0 84.7
RocketKV 100.0 100.0 100.0 82.0 90.0 70.0 94.0 98.0 89.6 76.4 86.7 78.0 56.0 86.2
Table 14: Results of RULER-24K for Mistral-7B-Ins-v0.2
@@
HSA 100.0 98.0 90.0 88.0 92.0 78.0 73.5 83.0 94.8 62.0 82.7 70.0 58.0 82.3
RocketKV 100.0 100.0 98.0 82.0 92.0 76.0 89.5 94.5 97.2 48.2 84.7 60.0 56.0 82.9
Table 15: Results of RULER-32K for Mistral-7B-Ins-v0.2
@@
HSA 20.8 29.2 43.2 33.5 24.2 14.1 31.0 22.9 26.5 66.0 84.2 12.9 28.5 55.3 59.3 0.0
RocketKV 20.3 29.2 42.8 33.5 24.9 14.5 30.5 23.3 26.5 66.0 84.2 12.8 28.5 53.5 56.9 0.0
Table 18: Results of RULER-8K for LongChat-7B-v1.5
@@
HSA 100.0 100.0 96.0 98.0 100.0 80.0 96.5 69.5 92.8 34.4 82.7 66.0 58.0 82.6
RocketKV 100.0 100.0 96.0 98.0 100.0 80.0 97.0 70.0 93.2 33.6 82.7 66.0 58.0 82.7
Table 19: Results of RULER-16K for LongChat-7B-v1.5
@@
HSA 100.0 100.0 96.0 92.0 98.0 52.0 83.0 53.0 79.2 26.4 70.7 62.0 44.0 73.6
RocketKV 100.0 100.0 98.0 92.0 100.0 46.0 83.0 54.0 78.8 24.6 72.0 62.0 44.0 73.4
Table 20: Results of RULER-24K for LongChat-7B-v1.5
@@
HSA 100.0 100.0 98.0 88.0 82.0 22.0 58.5 38.5 67.2 12.2 66.0 46.0 34.0 62.5
RocketKV 100.0 100.0 98.0 90.0 84.0 12.0 63.5 37.5 68.8 8.2 66.7 48.0 34.0 62.4
Table 21: Results of RULER-32K for LongChat-7B-v1.5